OpenStudy (anonymous):

preform the indicated operation. 3/x+5/3x^2 please help me!!!

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

You need a common denominator.

OpenStudy (anonymous):

except for the last demonitor.. its 3x^2 not x^2

OpenStudy (ivettef365):

I meant 3 5 ------- + ---------- x 3x^2

OpenStudy (mathstudent55):

What is the least common denominator for x and 3x^2?

OpenStudy (anonymous):

yes thats correct

OpenStudy (ivettef365):

ok, so what is the LCD ??

OpenStudy (anonymous):

i dont know?

OpenStudy (anonymous):

3?

OpenStudy (ivettef365):

actually it is 3x^2

OpenStudy (ivettef365):

then you have 3 5 3(3x) + 5 ----- + ------- = ------------------ x 3x^2 3x^2

OpenStudy (ivettef365):

x/3x^2 = 3x that is why you multiply the 3 (3x) and 3x^2/3x^2 = 1, so the 5(1) = 5

OpenStudy (ivettef365):

do you understand

OpenStudy (anonymous):

yes i get it. thank you

OpenStudy (ivettef365):

sorry the other way 3x^2/x = 3x

OpenStudy (anonymous):

so the answer is?

OpenStudy (ivettef365):

can you figure it out

OpenStudy (anonymous):

not really

OpenStudy (ivettef365):

ok is just a matter of adding 3(3x) = 9x so you have 9x +5 ------ 3x^2